# TinyLlama: An Open-Source Small Language Model

### v1

@@ 4 Conclusion
In this paper, we introduce TinyLlama, an open-source, small-scale language model. To promote transparency in the open-source LLM pre-training community, we have released all relevant information, including our pre-training code, all intermediate model checkpoints, and the details of our data processing steps. With its compact architecture and promising performance, TinyLlama can enable end-user applications on mobile devices, and serve as a lightweight platform for testing a wide range of innovative ideas related to language models. We will leverage the rich experience accumulated during the open, live phase of this project and aim to develop improved versions of TinyLlama, equipping it with a diverse array of capabilities to enhance its performance and versatility across various tasks. We will document further findings and detailed results in upcoming reports.
@@ Acknowledgements
We express our gratitude to the open-source community for their strong support during the open, live phase of our research. Special thanks go to Qian Liu, Longxu Dou, Hai Leong Chieu, and Larry Law for their help to our project. This research/project is supported by Ministry of Education, Singapore, under its Academic Research Fund (AcRF) Tier 2 Programme (MOE AcRF Tier 2 Award No.: MOE-T2EP20122-0011), Ministry of Education, Singapore, under its Tier 3 Programme (The Award No.: MOET320200004), the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Program (AISG Award No: AISG2-RP-2020-016), an AI Singapore PhD Scholarship (AISG Award No: AISG2-PhD-2021-08-007), an SUTD Kick-Starter Project (SKI 2021_03_11), and the grant RS-INSUR-00027-E0901-S00. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.

### v2

@@ 4 Conclusion
In this paper, we introduce TinyLlama, an open-source, small-scale language model. With its compact architecture and promising performance, TinyLlama can enable end-user applications on mobile devices, and serve as a lightweight platform for testing a wide range of innovative ideas related to language models. We also additionally verify the effectiveness of multi-stage training and data schedule through our v1.1 series. To promote transparency in the open-source language model pre-training community, we have released all relevant information, including our pre-training code, intermediate model checkpoints, and the details of our data processing steps. We hope that TinyLlama will make a valuable contribution to the community as we aim to democratize language model research and provide useful insights for future research in this field.
@@ Acknowledgements
We express our gratitude to the open-source community for their strong support during the open, live phase of our research. Special thanks go to Qian Liu, Longxu Dou, Hai Leong Chieu, and Larry Law for their help to our project. We would like to thank the National Supercomputing Centre (NSCC) Singapore for their support in training our model v1.1. This research/project is supported by Ministry of Education, Singapore, under its Academic Research Fund (AcRF) Tier 2 Programme (MOE AcRF Tier 2 Award No.: MOE-T2EP20122-0011), Ministry of Education, Singapore, under its Tier 3 Programme (The Award No.: MOET320200004), the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Program (AISG Award No: AISG2-RP-2020-016), an AI Singapore PhD Scholarship (AISG Award No: AISG2-PhD-2021-08-007), an SUTD Kick-Starter Project (SKI 2021_03_11), and the grant RS-INSUR-00027-E0901-S00. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.
Appendix A Data sampling ratio for TinyLlama v1.1
The data sampling ratio for TinyLlama v1.1 Math&Code and Chinese model are shown in Table 6 and 7.
Table 6: Data sampling ratio for TinyLlama-Math&Code
Basic pretraining Continual pretraining with specific domain Cooldown
Slimpajama 100.0% 75.0% 75.0%
Starcoder - 15.0% 15.0%
Proof pile - 10.0% 10.0%
Table 7: Data sampling ratio for TinyLlama-v1.1-Chinese
Basic pretraining Continual pretraining with specific domain Cooldown
Slimpajama 100.0% 50.0% 50.0%
Skypile - 50.0% 50.0%